OpenStudy (anonymous):

Passing through (-3,3) and parallel to the line whose equation is 3x-2y=5?

OpenStudy (nick67):

Rewriting your equation in the form:\[y = m x + n\] we have: \[y = 3/2 x - 5/2\]Then we calculate y value for x = -3, obtaining y = -7 ; to obtain the second line we need the same m factor (3/2) so that the two lines are parallel, but n factor should be with an offset of +10 (difference between 3 and -7) , i.e. :\[y = 3/2 x - 5/2 + 10 = 3/2 x + 15/2\]